### v1

# Power monoids and their arithmetic: a survey

Abstract
The non-empty finite subsets of a multiplicatively written monoid form a monoid in their own right, and so do the finite subsets that contain the identity element. Partly due to their unusual arithmetic properties, these structures, known as power monoids, have attracted increasing attention in recent years and have in turn stimulated growing interest in new perspectives in factorization theory, better suited to non-cancellative settings. We survey these developments and briefly review some related aspects.
Keywords: power monoids, factorizations, arithmetic properties of semigroups.
1 Introduction.
Let $S$ be a multiplicatively written semigroup (see Sect. 2 for notation and terminology). Endowed with the binary operation of setwise multiplication induced by $S$ on its power set and defined by
$XY:=\{xy:x\in X,\,y\in Y\},\qquad\text{for all }X,Y\subseteq S,$
the non-empty subsets of $S$ themselves form a semigroup, denoted by $\mathcal{P}(S)$. Furthermore, the family of all non-empty finite subsets of $S$ forms a subsemigroup of $\mathcal{P}(S)$, denoted by $\mathcal{P}_{\mathrm{fin}}(S)$. Obviously, $\mathcal{P}(S)=\mathcal{P}_{\mathrm{fin}}(S)$ if and only if $S$ is finite. We call $\mathcal{P}(S)$ the large power semigroup and $\mathcal{P}_{\mathrm{fin}}(S)$ the finitary power semigroup of $S$, and refer to either structure generically as a power semigroup.
If the semigroup $S$ is written additively, we adopt the same convention for any subsemigroup of $\mathcal{P}(S)$. This amounts, in practice, to the fact that the operation on $\mathcal{P}(S)$ maps a pair $(X,Y)$ of non-empty subsets of $S$ to their sumset
$X+Y:=\{x+y:x\in X,\;y\in Y\},$
as opposed to the setwise product of $X$ by $Y$ used in the multiplicative setting.
Power semigroups provide a natural algebraic framework for a variety of key problems in additive number theory and related areas [6, 47, 46, 19], including Sárközy’s conjecture on the “additive irreducibility” of the set of [non-zero] squares of a finite field of prime order [36, Conjecture 1.6] and Ostmann’s conjecture [29, p. 13] on the “asymptotic additive irreducibility” of the set of (positive rational) primes.
Their first explicit appearance can be traced back to a 1953 paper by Dubreil [15], where, however, the author argued in favor of including the empty set (see the unnumbered remark from loc. cit., p. 281). Their systematic investigation, begun in the late 1960s, was pursued intensively throughout the 1980s and 1990s.
By that time, researchers were mainly interested in the study of varieties (that is, classes of semigroups that are closed under taking homomorphic images, subsemigroups, and finite direct products) generated by $\mathcal{P}(S)$ as $S$ ranges over a specified family of finite semigroups. A catalyst of these early developments was the role of power semigroups in the study of formal languages and automata [3].
Power semigroups have also been central to the ongoing development of an arithmetic theory of semigroups and rings, aiming to extend the classical theory of factorizations [21] beyond its traditional boundaries. These aspects of the topic form the core of the present survey and will be discussed in much greater detail in Sections 4 and 5. In the remainder of this introduction, we will instead briefly review other directions of research that have contributed to the revival of the subject after it remained more or less dormant for roughly twenty years.
Let us just mention here that power semigroups are “strongly non-cancellative structures” (for instance, the large power semigroup of a finite group contains non-trivial idempotents, except in trivial cases), and the classical theory of factorizations is ill-suited to handle such situations. For this reason, in Section 3 we review the basics of an “extended theory of factorizations”, which provides tools to circumvent these and other obstacles inherent in the classical approach.
Isomorphism problems.
A turning point in the history of the topic was marked by a paper of Tamura and Shafer [40] that has eventually led to the following questions. Here and throughout, the symbol $\cong$ denotes isomorphism, and every morphism between semigroups is understood to be a semigroup homomorphism.
Questions 1.1.
Prove or disprove that, for all $H$ and $K$ in a given class of semigroups,
  1. (1)
    $\mathcal{P}(H)\cong\mathcal{P}(K)$ if and only if $H\cong K$.
  2. (2)
    $\mathcal{P}_{\mathrm{fin}}(H)\cong\mathcal{P}_{\mathrm{fin}}(K)$ if and only if $H\cong K$.
The crux of these questions lies in the “only if” direction. Indeed, if $f\colon H\to K$ is a semigroup isomorphism, then the map $f^{\ast}$ from $\mathcal{P}(H)$ to $\mathcal{P}(K)$ defined by
$f[X]:=\{f(x)\colon x\in X\},\qquad\text{for every non-empty }X\subseteq H,$ (1)
is itself an isomorphism, called the augmentation of $f$. Moreover, the restriction of $f^{\ast}$ to the non-empty finite subsets of $H$ yields an isomorphism from $\mathcal{P}_{\mathrm{fin}}(H)$ onto $\mathcal{P}_{\mathrm{fin}}(K)$, since clearly $f^{\ast}(X)\in\mathcal{P}_{\mathrm{fin}}(K)$ for all $X\in\mathcal{P}_{\mathrm{fin}}(H)$. See [44, Remark 4] and [19, Section 1] for further details and context.
The answer to Question 1.1(1) is negative for the class of all semigroups [28]; is positive for groups [37], semilattices [27, p. 218], Clifford semigroups [17, Theorem 4.7], cancellative commutative semigroups [44, Corollary 1], etc.; and is open for finite semigroups [24, p. 5], despite some authors having claimed the opposite based on results announced in [39] but never proved.
As for Question 1.1(2), very little seems to be known outside the case where $\mathscr{C}$ is a class of finite semigroups contained in any of the classes that are already addressed by the “positive results” reviewed in the previous paragraph (note that $\mathcal{P}(S)=\mathcal{P}_{\mathrm{fin}}(S)$ if and only if $S$ is finite). More precisely, Bienvenu and Geroldinger have shown in [6, Theorem 3.2(3)] that the question has an affirmative answer for numerical monoids, and the conclusion has been generalized to cancellative commutative semigroups in [44, Corollary 1].
Power monoids.
Investigations in the past few years have addressed some variants of Questions 1.1, inspired by the work of Bienvenu and Geroldinger in [6].
Specifically, let $M$ be a multiplicatively written monoid, with $1_{M}$ denoting its identity and $M^{\times}$ its unit group. $\mathcal{P}(M)$ itself is then a monoid, and $\mathcal{P}_{\mathrm{fin}}(M)$ is a submonoid of $\mathcal{P}(M)$, their identity being the singleton $\{1_{M}\}$. In addition, each of the families
$\mathcal{P}_{\mathrm{fin},\times}(M):=\{X\in\mathcal{P}_{\mathrm{fin}}(M):X\cap M^{\times}\neq\emptyset\}$
and
$\mathcal{P}_{\mathrm{fin},1}(M):=\{X\in\mathcal{P}_{\mathrm{fin}}(M):1_{M}\in X\}\subseteq\mathcal{P}_{\mathrm{fin},\times}(M),$
first considered by Fan and Tringali in [16], is a submonoid of $\mathcal{P}_{\mathrm{fin}}(M)$. Accordingly, $\mathcal{P}_{\mathrm{fin}}(M)$ is called the finitary power monoid, $\mathcal{P}_{\mathrm{fin},\times}(M)$ the restricted finitary power monoid, and $\mathcal{P}_{\mathrm{fin},1}(M)$ the reduced finitary power monoid of $M$. We will refer generically to any of these structures as a power monoid, and write $\mathcal{P}_{\mathrm{fin},0}(M)$ instead of $\mathcal{P}_{\mathrm{fin},1}(M)$ when $M$ is written additively.
Recent work on power monoids has mainly focused on $\mathcal{P}_{\mathrm{fin},1}(M)$. Most notably, Tringali and Yan [47] posed the following analogue of Questions 1.1, where, as before, the symbol $\cong$ denotes a semigroup isomorphism.
Question 1.2.
Prove or disprove that, for all $H$ and $K$ in a given class of monoids, $\mathcal{P}_{\mathrm{fin},1}(H)\cong\mathcal{P}_{\mathrm{fin},1}(K)$ if and only if $H\cong K$.
Once again, the core of Question 1.2 lies in proving the “only if ” direction. Indeed, if $f$ is a (semigroup) isomorphism from a monoid $H$ to a monoid $K$, then $f$ maps the identity of $H$ to the identity of $K$ (see, for instance, the last lines of [44, Sec. 2]). Since $f[X]$ is finite for every non-empty finite $X\subseteq H$, it is thus clear that the augmentation of $f$, as defined by Eq. (1), restricts to an isomorphism from $\mathcal{P}_{\mathrm{fin},1}(H)$ to $\mathcal{P}_{\mathrm{fin},1}(K)$.
With these preliminaries in mind, it was shown in [47] that Question 1.2 has a negative answer for left- or right-cancellative monoids (loc. cit., Examples 1.2), and a positive answer for (rational) Puiseux monoids (loc. cit., Theorem 2.5), that is, submonoids of the non-negative rational numbers under addition. The latter result confirmed the “Bienvenu–Geroldinger conjecture”, namely, Conjecture 4.7 in [6].
Rago [32] subsequently proved that Question 1.2 has a negative answer within the class of cancellative commutative monoids. More precisely, he showed that if $H$ and $K$ are commutative valuation monoids with trivial unit groups and isomorphic quotient groups, then $\mathcal{P}_{\mathrm{fin},1}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin},1}(K)$. More recently, he characterized all pairs $(H,K)$ of non-isomorphic cancellative commutative monoids such that $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ are isomorphic [30, Theorem 18]. In particular, he obtained in [30, Theorem 14] that if $H$ and $K$ have non-trivial unit groups and $\mathcal{P}_{\mathrm{fin},1}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin},1}(K)$, then $H$ is isomorphic to $K$.
Meanwhile, Tringali and Yan have established that Question 1.2 has an affirmative answer for torsion groups [48], and Rago (private communication) has recently announced a proof that the same conclusion extends to arbitrary groups.
Symmetry and rigidity.
The study of automorphisms across a wide range of objects and categories (from graphs to fields, from topological spaces to Banach algebras) is closely connected to many other properties of interest and often provides a deeper understanding of various aspects of their theory.
With these thoughts in mind, we gather from the previous subsection that the augmentation of an automorphism $f$ of a semigroup $S$ restricts to an automorphism of the finitary power semigroup $\mathcal{P}_{\mathrm{fin}}(S)$. This gives a canonical embedding $\Phi\colon\mathrm{Aut}(S)\hookrightarrow\mathrm{Aut}(\mathcal{P}_{\mathrm{fin}}(S))$, and it is natural to ask whether $\Phi$ is a (group) isomorphism. Mutatis mutandis, the same question arises for $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},1}(S))$ when $S$ is a monoid.
In the aftermath of their proof [47, Theorem 2.5] of [6, Conjecture 4.7], Tringali and Yan obtained in [46, Theorem 3.2] that the automorphism group of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is cyclic of order two, and conjectured [46, Sec. 4] that $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},0}(H))$ is trivial for every numerical monoid $H\neq\mathbb{N}$. As a follow-up, Tringali and Wen [45, Theorem 3.5] have shown that the automorphism group of the finitary power monoid of $\mathbb{Z}$ (the additive group of integers) is isomorphic to the direct product of a cyclic group of order two and the infinite dihedral group, in stark contrast with the fact that the only non-trivial automorphism of the same group is the map $x\mapsto-x$. In a remarkable feat, Rago [31, Theorem 13] has subsequently proved that, for a finite abelian group $G$, the automorphism group of $\mathcal{P}_{\mathrm{fin},1}(G)$ is (canonically) isomorphic to $\mathrm{Aut}(G)$, except when $G$ is the Klein four-group, in which case $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},1}(G))$ is isomorphic to the direct product of two copies of the symmetric group of degree three [31, Example 2]. Even more recently, Wong et al. [49, Theorem 1.3] have established that the automorphism group of the finitary power semigroup $\mathcal{P}_{\mathrm{fin}}(S)$ of a numerical semigroup $S$ is trivial unless $S=\llbracket k,\infty\rrbracket$ for some integer $k\geq 0$, in which case $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin}}(S))$ is cyclic of order two.
2 Formalities.
We denote by $\mathbb{N}$ the additive monoid of non-negative integers, by $\mathbb{N}^{+}$ the set of positive integers, by $\mathbb{Z}$ the additive group of integers, and by $|X|$ the cardinality of a set $X$. Unless otherwise stated, we reserve the letters $m$ and $n$ (with or without subscripts) for positive integers. Given $a,b\in\mathbb{N}\cup\{\infty\}$, we let $\llbracket a,b\rrbracket:=\{x\in\allowbreak\mathbb{N}\colon\allowbreak a\leq x\leq b\}$ be the (discrete) interval from $a$ to $b$.
If not explicitly specified, we write all semigroups (and monoids) multiplicatively. An element $a$ in a semigroup $S$ is cancellative if $ax\neq ay$ and $xa\neq ya$ for all $x,y\in S$ with $x\neq y$; the semigroup itself is cancellative if each of its elements is.
A unit of a monoid $M$ with identity (element) $1_{M}$ is an element $u\in M$ for which there exists a (provably unique) element $v\in M$, accordingly denoted by $u^{-1}$ and called the inverse of $u$ (in $M$), with the property that $uv=vu=1_{M}$. We assume that every submonoid of a monoid is unitary, meaning that it has the same identity.
The monoid $M$ is Dedekind-finite if, for any $x,y\in M$, the equality $xy=1_{M}$ implies $yx=1_{M}$; unit-cancellative if $xy\neq x\neq yx$ for all $x,y\in M$ with $y$ a non-unit; acyclic if $uxv\neq x$ for all $u,v,x\in M$ whenever of $u$ or $v$ is a non-unit; and torsion-free if the only finite-order element of $M$ is the identity. Here and throughout, the order of an element $x\in M$ is the size of the set $\{x^{k}:\allowbreak k\in\allowbreak\mathbb{N}\}$ if this set is finite; otherwise, the order is $\infty$.
To the best of our knowledge, unit-cancellativity was first studied in the commutative setting by Rosales et al. [35, Section 4, p. 72] (under a different name), and in the non-commutative setting by Fan and Tringali [16, Section 2.1]. Acyclicity, on the other hand, was first introduced in [42, Definition 4.2]. Commutative or unit-cancellative monoids are Dedekind-finite [42, Proposition 4.3(i)]. Moreover, every acyclic monoid is unit-cancellative, and the converse holds in the commutative setting.
Given a set $A$, we denote by $\mathscr{F}^{+}(A)$ the free semigroup over $A$ and write $\ast_{A}$ for its operation; when no ambiguity is likely to arise, we drop the subscript $A$ from this notation. As a set, $\mathscr{F}^{+}(A)$ consists of all non-empty finite tuples of elements of $A$, and $\ast_{A}$ is the operation of concatenation of such tuples.
We refer to an element of $\mathscr{F}^{+}(A)$ as an $A$-word, or simply a word when there is no serious risk of confusion. If $\mathfrak{u}$ is an $A$-word, its length is the unique integer $k\geq 1$ such that $\mathfrak{u}$ is a $k$-tuple of elements of $A$. In this case, $\mathfrak{u}=u_{1}\ast\cdots\ast u_{k}$ for some uniquely determined $u_{1},\ldots,u_{k}\in A$, called the letters of $\mathfrak{u}$.
Further notation and terminology, if not explained when first used, are standard or should be clear from the context. In particular, we address the reader to Howie’s monograph [25] for basic aspects of semigroup theory.
3 New perspectives on an old classic.
Throughout this section, $M$ denotes a monoid, with no further assumptions unless otherwise specified.
Roughly speaking, the theory of factorizations is the study of a broad spectrum of phenomena arising from the possibility or impossibility of extending the Fundamental Theorem of Arithmetic from the integers to more abstract settings where a set comes endowed with a binary operation. The theory has traditionally focused on domains and cancellative monoids, primarily in the commutative setting [21, 20, 22, 4, 38], and the building blocks used in factorizations are atoms, that is, non-units that cannot be written as a product of two other non-units. Only in the past few years have researchers taken the first steps toward a systematic study of factorizations in non-cancellative (and non-commutative) settings, using building blocks that are not necessarily atoms; see [5, Section 2.4] for a review of some earlier scattered results along these lines.
Power monoids are among the driving forces behind these developments: they exhibit a rich arithmetic, which makes them an ideal testing ground for some key aspects of an “extended theory of factorizations” whose foundations were laid down by Tringali et al. in [16, 5, 42, 43, 11, 10, 12]. See [2, 8, 7, 18] for applications to Boolean lattices, monoids of ideals, and semilattices of numerical monoids.
A key feature of these new perspectives is the idea of pairing the monoid $M$ with a preorder (that is, a reflexive and transitive binary relation on $M$) and letting the preorder induce on $M$ a relative notion of irreducibility [42, Section 3]. In this way, the theory gains tremendous flexibility, as neatly explained by Cossu in [9], allowing, for instance, the study of factorizations in non-trivial contexts such as groups, where the classical theory is inapplicable because there are no atoms, or in rings with non-zero zero divisors, where the classical theory loses significance as its invariants tend to blow up in a more or less predictable way.
Another feature of this extended theory, specifically designed to counter, or at least mitigate, the aforementioned blow-up phenomena, is the idea of imposing a “minimality condition” on factorizations, an idea first introduced in [5, Section 4] within the restricted framework of the classical theory and further developed in [10, 12, 13].
Below, we briefly review the basics of this generalized approach that are most relevant to the study of power monoids in the subsequent sections.
Irreducibles vs atoms.
To avoid unnecessary technicalities, we shall not pursue the broader generality alluded to above, but restrict our attention to the divisibility preorder $\mid_{M}$, that is, the binary relation on $M$ defined, for all $x,y\in M$, by
$x\mid_{M}y\quad\text{if and only if}\quad y\in MxM.$
If $x\mid_{M}y$, we say that $x$ is a divisor of $y$ (in $M$), or $x$ divides $y$, or $y$ is divisible by $x$. If $x\mid_{M}y$ and $y\mid_{M}x$, we write $x\simeq_{M}y$. If $x$ does not divide $y$, we write $x\nmid_{M}y$. We call $x$ a unit-divisor if $x$ is a divisor of the identity $1_{M}$, and a proper divisor of $y$ if $x\mid_{M}y$ but $y\nmid_{M}x$. For later reference, it is convenient to set
$M^{\sharp}:=\{x\in M:x\nmid_{M}1_{M}\}.$
A submonoid $N$ of $M$ is divisor-closed if every divisor in $M$ of an element $a\in N$ is itself an element of $N$; in this case, we may equally say that $N$ is divisor-closed in $M$. Note that every unit divides every element of $M$, and $M$ is Dedekind-finite if and only if $M^{\sharp}=M\setminus M^{\times}$ (that is, every unit-divisor is a unit).
Within this framework, a key role is played by the $\mid_{M}$-irreducibles, which we simply call the irreducibles, or irreducible elements, of $M$. These are the elements $a\in M^{\sharp}$ that admit no factorization $a=bc$ with $b$ and $c$ proper divisors of $a$ in $M$ that are not unit-divisors. In particular, it follows from the previous observations that, in a Dedekind-finite monoid, an irreducible is a non-unit that does not factor as a product of two proper divisors. In addition to irreducibles, we shall occasionally encounter quarks, that is, elements whose only proper divisors divide the identity, yet which are not themselves unit-divisors.
Atoms (in the sense of the classical theory) and quarks are irreducibles, but the converse need not be true. For instance, the zero element $0_{R}$ of a domain $R$ is irreducible in the multiplicative monoid of $R$, yet it is not an atom, since it is idempotent. Moreover, if $R$ is not a skew field, then $0_{R}$ is not a quark either, because every non-unit $x\in R$ divides $0_{R}$. For a more interesting example, see Theorem 4.7(ii).
Proposition 3.1.
Every atom of a monoid is a quark.
Proof.
Let $a$ be an atom of $M$ (if $M$ has no atoms, the statement is vacuously true), and let $b\in M$ be a divisor of $a$ but not a unit-divisor. By [16, Lemma 2.2(i)], $M$ is Dedekind-finite, and hence $b$ is a non-unit. We need to show that $a\mid_{M}b$.
Since $b\mid_{M}a$, there exist $u,v\in M$ such that $a=ubv$; and since $a$ is an atom, this can only happen if each of the sets $\{u,bv\}$ and $\{ub,v\}$ contains a unit. By the Dedekind-finiteness of $M$ and the fact that $b\notin M^{\times}$, it follows that both $u$ and $v$ are units. Hence $b=u^{-1}av^{-1}$, that is, $a$ divides $b$ (as desired). ∎
For the reader’s convenience, we summarize the mutual relationships between irreducibles, atoms, and quarks in the following diagram:
$\{\text{atoms}\}\subseteq\{\text{quarks}\}\subseteq\{\text{irreducibles}\}.$
It should be noted that our distinction between irreducibles and atoms is a delicate point and a common source of confusion. While in the classical theory the two terms are used interchangeably, the former notion is considerably more flexible than the latter and better suited to the extended version of the theory that we consider here; see, e.g., [42, Theorem 3.10], [11, Theorem 5.19], and [10, Theorem 5.1].
Factorizations and lengths.
The monoid $M$ is factorable (resp., atomic) if every $x\in M^{\sharp}$ can be written as a product $a_{1}\cdots a_{n}$ of irreducibles (resp., atoms), in which case the word $a_{1}\ast\allowbreak\cdots\ast a_{m}$ (an element of the free semigroup over $M$) is called a factorization (resp., an atomic factorization) of $x$ (over $M$) and $m$ is called a length (resp., an atomic length) of $x$. We denote by $\mathscr{Z}_{M}(x)$ the set of factorizations of $x$, and by $\mathsf{L}_{M}(x)$ the set of lengths of $x$.
A factorization $\mathfrak{a}\in\mathscr{Z}_{M}(x)$ is equivalent to a factorization $\mathfrak{b}\in\mathscr{Z}_{M}(x)$ if the length $m$ of $\mathfrak{a}$ is equal to the length $n$ of $\mathfrak{b}$ and there exists a permutation $\sigma$ of the interval $\llbracket 1,m\rrbracket$ such that $a_{i}\simeq_{M}b_{\sigma(i)}$ for every $i\in\llbracket 1,m\rrbracket$, where $a_{i}$ and $b_{\sigma(i)}$ are the $i^{\text{th}}$ letter of $\mathfrak{a}$ and the $\sigma(i)^{\text{th}}$ letter of $\mathfrak{b}$. Of course, “being equivalent” is an equivalence relation on $\mathscr{Z}_{M}(x)$. In addition, $\mathfrak{a}$ is a minimal factorization (of $x$) if there do not exist a factorization $\alpha_{1}\ast\cdots\ast\alpha_{k}\in\mathscr{Z}_{M}(x)$ of length $k$ (strictly) smaller than $m$ and an injective map $\tau\colon\llbracket 1,k\rrbracket\to\llbracket 1,m\rrbracket$ such that $\alpha_{i}\simeq_{M}a_{\tau(i)}$ for each $i\in\llbracket 1,k\rrbracket$, in which case $m$ is called a minimal length of $x$. We denote by $\mathsf{L}_{M}^{\sf m}(x)$ the set of minimal lengths and by $\mathscr{Z}_{M}^{\sf m}(x)$ the set of minimal factorizations of $x$. It is fairly obvious that $x$ has a factorization if and only if it has a minimal factorization; moreover,
$\mathsf{L}_{M}^{\sf m}(x)\subseteq\mathsf{L}_{M}(x)\quad\text{and}\quad\mathscr{Z}_{M}^{\sf m}(x)\subseteq\mathscr{Z}_{M}(x).$
For convenience, we extend some of the previous definitions, by setting $\mathsf{L}_{M}(1_{M})=\mathsf{L}_{M}^{\sf m}(1_{M}):=\{0\}\subseteq\mathbb{N}$ and $\mathsf{L}_{M}(u)=\mathsf{L}_{M}^{\sf m}(u):=\emptyset$ for every non-identity unit-divisor $u\in M$. (Note that a non-empty product of irreducibles cannot be a unit-divisor.) Accordingly, it is natural to consider the families
$\mathscr{L}(M):=\{\mathsf{L}_{M}(y):y\in M\}\setminus\{\emptyset\}$
and
$\mathscr{L}^{\sf m}(M):=\{\mathsf{L}_{M}^{\sf m}(y):y\in M\}\setminus\{\emptyset\}.$
These are subfamilies of the power set of $\mathbb{N}$, and both contain the singleton $\{0\}$. Moreover, our definitions ensure that if $0\in\mathsf{L}_{M}(y)$ for some $y\in M$, then $y$ is the identity $1_{M}$, and hence $\mathsf{L}_{M}(y)=\mathsf{L}_{M}^{\sf m}(y)=\{0\}$.
Each $L\in\mathscr{L}(M)$ is called a length set of $M$, and $\mathscr{L}(M)$ the system of lengths of the monoid. In a similar vein, each $L\in\mathscr{L}^{\sf m}(M)$ is called a minimal length set of $M$, and $\mathscr{L}^{\sf m}(M)$ the system of minimal lengths.
A fundamental tenet of factorization theory is that, apart from trivial cases, the arithmetic of a monoid $M$ is controlled by the arithmetic of its proper divisor-closed submonoids. From [10, Corollary 3.6], we know that if $N$ is a divisor-closed submonoid of $M$, then $N$ and $M$ share the same unit-divisors, and every irreducible (resp., atom) of $N$ is also an irreducible (resp., atom) of $M$. We are thus led to the following proposition, which is a (very) special case of [10, Proposition 3.5].
Proposition 3.2.
If $N$ is a divisor-closed submonoid of the monoid $M$, then every [minimal] factorization of an element $x\in N$ is also a [minimal] factorization of $x$ in $M$. In particular, the system of [minimal] lengths of $N$ is contained in that of $M$.
Length sets are among the most studied arithmetic invariants in the classical theory of factorizations. In this regard, it is worth noting that, for an acyclic monoid, every irreducible is an atom [42, Corollary 4.4], so there is no distinction between factorizations and atomic factorizations.
We say that $M$ is factorial (shortly, UF) if it is factorable and any two factorizations of an element in $M^{\sharp}$ are equivalent; finite-factorial (shortly, FF) if it is factorable and, for each $x\in M^{\sharp}$, the set of factorizations of $x$ is finite upon identifying equivalent factorizations; half-factorial (shortly, HF) if it is factorable and every length set is a singleton; and bounded-factorial (shortly, BF) if it is factorable and every length set is finite.
These definitions are modeled on their classical counterparts [21, Theorem 1.1.10.2 and Definitions 1.1.1.4, 1.2.6, 1.3.1, and 1.5.1], and it is fairly obvious that the implications shown in the following diagram hold:
HFUFBFfactorableFF
Analogous definitions can be formulated for minimal factorizations. More explicitly, $M$ is UmF if it is factorable and any two minimal factorizations of a non-unit are equivalent; FmF if, for each non-unit $x\in M$, the set of minimal factorizations of $x$, up to equivalence, is finite and non-empty; HmF if it is factorable and any two minimal factorizations of the same non-unit have the same length; and BmF if it is factorable and every set of minimal lengths is finite.
Lastly, restricting to atomic factorizations yields the corresponding notions of UF-atomic, …, BF-atomic, UmF-atomic, …, and BmF-atomic monoid.
4 Arithmetic in power monoids.
Throughout, $H$ denotes a monoid, and no additional assumptions on $H$ are made unless explicitly stated.
In this section, we survey several results concerning irreducible, atomic, and minimal factorizations in the reduced finitary power monoid $\mathcal{P}_{\mathrm{fin},1}(H)$ of $H$. Related structures, including $\mathcal{P}_{\mathrm{fin},\times}(H)$ and $\mathcal{P}_{\mathrm{fin}}(H)$, will be mentioned only occasionally, since the current state of knowledge does not yet allow for a systematic treatment comparable to that of $\mathcal{P}_{\mathrm{fin},1}(H)$. Some partial results, notably by Fan and Tringali [16], Antoniou and Tringali [5], Bienvenu and Geroldinger [6], Gonzalez et al. [23], and Gotti et al. [1, 14], will not be discussed here.
Overture.
There are many arguments supporting the claim that $\mathcal{P}_{\mathrm{fin},1}(H)$ is, in a sense, tamer than other power monoids. One basic point to keep in mind is that
$\text{if }X\text{ divides }Y\text{ in }\mathcal{P}_{\mathrm{fin},1}(H),\text{ then }X\subseteq Y.$ (2)
Another elementary observation (see, e.g., Proposition 3.2(ii) in [5]) is that
$|XY|\geq\max(|X|,|Y|),\qquad\text{for all }X,Y\in\mathcal{P}_{\mathrm{fin},\times}(H).$ (3)
Eqs. (2) and (3) enter into the proof of the next two results, where a monoid $M$ is said to satisfy the ascending chain condition (for short, ACC) on principal two-sided ideals if there is no sequence $x_{1},\allowbreak x_{2},\dots$ in $M$ such that $Mx_{n}M\subsetneq Mx_{n+1}M$ (or equivalently, $x_{n+1}$ is a proper divisor of $x_{n}$) for all $n$.
Theorem 4.1.
The following hold:
  1. (i)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is a Dedekind-finite monoid with trivial unit group, satisfies the ACC on principal two-sided ideals, and is factorable.
  2. (ii)
    $\mathcal{P}_{\mathrm{fin},1}(K)$ is a divisor-closed submonoid of $\mathcal{P}_{\mathrm{fin},1}(H)$ for every submonoid $K$ of $H$ (regardless of whether $K$ is divisor-closed in $H$).
Proof.
See [42, Proposition 4.11(i)–(ii)] for (i), and [5, Proposition 3.2(iii)] for (ii). ∎
Theorem 4.2.
The following hold:
  1. (i)
    The units of $\mathcal{P}_{\mathrm{fin},\times}(H)$ and $\mathcal{P}_{\mathrm{fin}}(H)$ are the singletons $\{u\}$ with $u\in H^{\times}$.
  2. (ii)
    $\mathcal{P}_{\mathrm{fin},\times}(H)$ is divisor-closed in $\mathcal{P}_{\mathrm{fin}}(H)$ if and only if $H$ is Dedekind-finite.
  3. (iii)
    $\mathcal{P}_{\mathrm{fin},\times}(H)$ is Dedekind-finite (regardless of any conditions on $H$), whereas $\mathcal{P}_{\mathrm{fin}}(H)$ is Dedekind-finite if and only if $H$ is.
  4. (iv)
    $\mathcal{P}_{\mathrm{fin},\times}(H)$ satisfies the ACC on principal two-sided ideals and is factorable.
Proof.
For (i), see [5, Proposition 3.2(iv)]. For (ii), use that $H$ embeds into $\mathcal{P}_{\mathrm{fin}}(H)$ via the homomorphism $x\mapsto\{x\}$, and recall that a monoid is Dedekind-finite if and only if a non-empty product is a unit precisely when each factor is a unit.
(iii) The Dedekind-finiteness of $\mathcal{P}_{\mathrm{fin},\times}(H)$ follows directly from (i) and Eq. (3). Moreover, for $\mathcal{P}_{\mathrm{fin}}(H)$ to be Dedekind-finite it is clearly necessary that $H$ is, since $H$ embeds into $\mathcal{P}_{\mathrm{fin}}(H)$ via the homomorphism $x\mapsto\{x\}$.
Now to finish the proof assume that $H$ is Dedekind-finite. If $XY=\{1_{H}\}$ for some $X,Y\in\mathcal{P}_{\mathrm{fin}}(H)$, then each of $X$ and $Y$ contains a unit of $H$, namely, $X,Y\in\mathcal{P}_{\mathrm{fin},\times}(H)$. This implies from the previous part that $X$ and $Y$ are units of $\mathcal{P}_{\mathrm{fin}}(H)$, which shows in turn that $\mathcal{P}_{\mathrm{fin}}(H)$ is Dedekind-finite.
(iv) Suppose, for a contradiction, that $\mathcal{P}_{\mathrm{fin},\times}(H)$ does not satisfy the ascending chain condition on principal two-sided ideals (ACCP, for short). There then exists a sequence $X_{1},X_{2},\dots$ in $\mathcal{P}_{\mathrm{fin},\times}(H)$ such that $X_{n+1}$ is a proper divisor of $X_{n}$ for each $n$. In particular, there exist $u_{n},v_{n}\in H^{\times}$ with $u_{n}X_{n+1}v_{n}\subseteq X_{n}$, and this inclusion must be strict; otherwise, $X_{n+1}=u_{n}^{-1}X_{n}v_{n}^{-1}$, making $X_{n}$ a divisor of $X_{n+1}$. It follows that $|X_{n+1}|=|u_{n}X_{n+1}v_{n}|<|X_{n}|$ for each $n$, which is impossible, as it would result in a strictly decreasing (infinite) sequence of positive integers.
Thus, $\mathcal{P}_{\mathrm{fin},\times}(H)$ satisfies the ACCP, which completes the proof, since $\mathcal{P}_{\mathrm{fin},\times}(H)$ is Dedekind-finite by (iii) and every Dedekind-finite monoid satisfying the ACCP is known to be factorable [42, Theorem 3.10]. ∎
Interlude on cancellation.
We have already mentioned that power monoids are far from being cancellative, which is one reason why much of the classical machinery of factorization theory breaks down when applied to the study of their arithmetic. The next results make this statement slightly more precise.
Proposition 4.3.
The following conditions are equivalent:
  1. (a)
    $H$ is trivial.
  2. (b)
    $\mathcal{P}_{\mathrm{fin}}(H)$ is cancellative.
  3. (c)
    $\mathcal{P}_{\mathrm{fin},\times}(H)$ is cancellative.
  4. (d)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is cancellative.
Proof.
Since every subsemigroup of a cancellative semigroup is cancellative, it is obvious that (a) $\Rightarrow$ (b) $\Rightarrow$ (c) $\Rightarrow$ (d). It remains to see that (d) $\Rightarrow$ (a).
If $H$ contains an element $x$ of order larger than $2$, then $\{1_{H},x\}^{2}\neq\{1_{H},\allowbreak x^{2}\}$, yet $\{1_{H},x\}^{3}=\{1_{H},x\}\{1_{H},x^{2}\}$. If, on the other hand, $H$ has an element $y$ of order $2$, then $\{1_{H}\}\neq\{1_{H},y\}=\{1_{H},y\}^{2}$. In either case, $\mathcal{P}_{\mathrm{fin},1}(H)$ is not cancellative. ∎
Although cancellativity fails dramatically in power monoids, it is still possible to guarantee certain “surrogates” (such as unit-cancellativity or torsion-freeness) under assumptions that, while strong, hold in many situations of interest.
Proposition 4.4.
If $H$ is a submonoid of a torsion-free group, then $\mathcal{P}_{\mathrm{fin},1}(H)$ is acyclic and, hence, unit-cancellative and torsion-free.
Proof.
Obviously, every acyclic monoid is unit-cancellative. Moreover, it is clear that any acyclic monoid $M$ with trivial unit group is torsion-free: if $x^{m}=x^{n}$ for some $x\in M$ and $m,n\in\mathbb{N}^{+}$ with $m<n$, then $x$ must be a unit and hence the identity. It is therefore sufficient to show that $\mathcal{P}_{\mathrm{fin},1}(H)$ is acyclic, as we know from Theorem 4.1(i) that the only unit of $\mathcal{P}_{\mathrm{fin},1}(H)$ is the identity.
Suppose that $AXB=X$ for some $A,B,X\in\mathcal{P}_{\mathrm{fin},1}(H)$. Since $H$ is contained in a torsion-free group, it follows from Kemperman’s inequality (see the unnumbered corollary to Theorem 5 on p. 251 of [26]) that
$|X|=|AXB|\geq|A|+|X|+|B|-2.$
Consequently, we must have $|A|+|B|\leq 2$, which is only possible if $|A|=|B|=1$, that is, $A=B=\{1_{H}\}$. This shows that $\mathcal{P}_{\mathrm{fin},1}(H)$ is acyclic (as desired). ∎
For the next proposition, we recall that a semigroup $S$ is linearly orderable if there exists a total order $\preceq$ on $S$ such that $x\prec y$ (that is, $x\preceq y$ and $x\neq y$) implies $uxv\prec uyv$ for all $u,v\in S$. For instance, every cancellative commutative semigroup is linearly orderable, and many more examples are listed in [41].
We also say that a monoid $M$ is strongly unit-cancellative if $xy\neq x$ and $yx\neq x$ for all $x,y\in M$ with $y\neq 1_{M}$. Of course, every strongly unit-cancellative monoid is unit-cancellative, and the converse holds when the unit group is trivial.
Proposition 4.5.
If $H$ is a linearly orderable monoid, then $\mathcal{P}_{\mathrm{fin}}(H)$ and $\mathcal{P}_{\mathrm{fin},\times}(H)$ are strongly unit-cancellative and torsion-free.
Proof.
For a proof that $\mathcal{P}_{\mathrm{fin}}(H)$ is strongly unit-cancellative, see Proposition 3.5(ii) in [16]. Since the properties of being strongly unit-cancellative or torsion-free pass to submonoids, it remains to show that $\mathcal{P}_{\mathrm{fin}}(H)$ is torsion-free.
To this end, let $\preceq$ be a total order on $H$ such that, whenever $x\prec y$, one has $uxv\prec uyv$ for all $u,v\in H$. Every non-empty finite subset $Y$ of $H$ has a $\preceq$-minimum and a $\preceq$-maximum; that is, there exist $y^{-},y^{+}\in Y$ such that $y^{-}\preceq y\preceq y^{+}$ for all $y\in Y$.
Next, fix a set $X\in\mathcal{P}_{\mathrm{fin}}(H)$ with $X\neq\{1_{H}\}$. It is clear that either $\min X\prec 1_{H}$ or $1_{H}\prec\max X$. We may assume without loss of generality that the latter holds (the other case being symmetric). If $m<n$, then
$\max X^{m}=m\max X\prec n\max X=\max X^{n},$
and hence $X^{m}\neq X^{n}$. It follows that $\mathcal{P}_{\mathrm{fin}}(H)$ is torsion-free (as desired). ∎
Minuetto of irreducibles.
We now turn to clarifying the relationship between irreducibles, atoms, and quarks. We begin with a key property of two-element sets.
Proposition 4.6.
$\{1_{H},x\}$ is a quark of $\mathcal{P}_{\mathrm{fin},1}(H)$ for each non-identity $x\in H$, while $\{u,x\}$ is a quark of $\mathcal{P}_{\mathrm{fin},\times}(H)$ for all $u\in H^{\times}$ and $x\in H$ with $u\neq x$.
Proof.
The first part is straightforward from Eq. (2), so we focus on the second.
Let $A\in\mathcal{P}_{\mathrm{fin},\times}(H)$ be a two-element set, and suppose that $X$ is a divisor of $A$ in $\mathcal{P}_{\mathrm{fin},\times}(H)$ but not a unit. By Proposition 4.2(i), $X$ has at least two elements, and there exist $S,T\in\mathcal{P}_{\mathrm{fin},\times}(H)$ such that $A=SXT$. Pick a unit $u\in S$ and a unit $v\in T$. Since $uXv\subseteq A$ and $2\leq|X|=|uXv|=|A|=2$, it follows that $A=uXv$, and hence $X=\allowbreak u^{-1}Av^{-1}$. Thus, AA divides $X$, thereby proving that $A$ is a quark. ∎
We continue with a characterization of quarks and atomicity in $\mathcal{P}_{\mathrm{fin},1}(H)$.
Theorem 4.7.
The following hold in $\mathcal{P}_{\mathrm{fin},1}(H)$:
  1. (i)
    Every irreducible is a quark.
  2. (ii)
    Every irreducible is an atom, with the exception of the two-element sets $\{1_{H},x\}\subseteq H$ for which $x^{2}=1_{H}$ or $x^{2}=x$.
Proof.
For (i), see [42, Proposition 4.11(iii)]. For (ii), see [13, Proposition 2.3], where it is however mistakenly stated (the proof is correct) that $X$ is irreducible but not an atom “if and only if $X=\{1_{H},x\}$ for some $x\in H$ such that $x^{2}=1_{H}$ or $x^{2}=x$” (one should add that $x\neq 1_{H}$). ∎
Corollary 4.8.
$\mathcal{P}_{\mathrm{fin},1}(H)$ is atomic if and only if $H$ has no unit of order two and its only idempotent is the identity $1_{H}$.
Proof.
Apply Theorems 4.1(i) and 4.7(ii), upon observing that, for every non-identity $x\in H$, the set $\{1_{H},x\}$ factors into a product of atoms if and only if it is an atom. ∎
Now we look for analogous characterizations in $\mathcal{P}_{\mathrm{fin},\times}(H)$.
Lemma 4.9.
Assume $H$ is Dedekind-finite, and let $A,X,Y\in\mathcal{P}_{\mathrm{fin},\times}(H)$ satisfy $A=XAY$. If $|A|\geq 3$ and $\max(|X|,|Y|)\geq 2$, then $A$ is not irreducible in $\mathcal{P}_{\mathrm{fin},\times}(H)$.
Proof.
Assume without loss of generality that $|X|\geq 2$ (the case $|Y|\geq 2$ is symmetric). Pick a unit $a\in A$. By the hypothesis that $A=XAY$, there exist $u\in X$, $b\in A$, and $v\in Y$ such that $a=ubv$, which, by the Dedekind-finiteness of $H$, is only possible if $u,v\in H^{\times}$. Since $uAv\subseteq A$ and $|uAv|=|A|$, it follows that $A=uAv$.
Now, using that $|X|\geq 2$, pick an element $x\in X$ with $x\neq u$. We gather from the above that $xav\in XAY=A=uAv$. Therefore, $xav=ua^{\prime}v$ for some $a^{\prime}\in A$, and hence $xa=ua^{\prime}$ (because $v$ is a unit and units are cancellative elements). It is then clear that $a\neq a^{\prime}$; otherwise, $a^{\prime}$ is a unit, and $xa=ua^{\prime}$ yields $x=u$ (a contradiction).
Set $B:=A\setminus\{a^{\prime}\}$. From the preceding considerations, it is immediate that
$\begin{split}A&=(A\setminus\{ua^{\prime}v\})\cup\{ua^{\prime}v\}=(uAv\setminus\{ua^{\prime}v\})\cup\{xav\}\\ &\subseteq uBv\cup xBv=\{u,x\}Bv\subseteq XAY=A.\end{split}$
Hence $A=\{u,x\}Bv$, which implies that $A$ is not irreducible in $\mathcal{P}_{\mathrm{fin},\times}(H)$.
Indeed, both $\{u,x\}$ and $Bv$ contain a unit of $H$ (in particular, $av\in Bv\cap H^{\times}$) and are therefore divisors of $A$ in $\mathcal{P}_{\mathrm{fin},\times}(H)$. Yet, by Eq. (3), neither $\{u,x\}$ nor $Bv$ is divisible by $A$, since $|\{u,x\}|=2<|A|$ and |Bv|=|B|=|A|1|Bv|=|B|=|A|-1. ∎
Proposition 4.10.
If $H$ is Dedekind-finite, then every irreducible of $\mathcal{P}_{\mathrm{fin},\times}(H)$ with at least two elements is an atom.
Proof.
Assume $A=XY$ for some $X,Y\in\mathcal{P}_{\mathrm{fin},\times}(H)$. We need to show that $X$ or $Y$ is a unit. Since $A$ is irreducible, either $X$ or $Y$ is divisible by $A$. Suppose we are in the former case (the other being analogous). There then exist $S,T\in\mathcal{P}_{\mathrm{fin},\times}(H)$ such that $X=SAT$ and hence $A=XY=SATY$. By Lemma 4.9, this is only possible if $|S|=|TY|=1$, which, by Eq. (3), yields $|Y|=1$ and hence proves, by Theorem 4.2(i), that $Y$ is a unit (as desired). ∎
Proposition 4.11.
Let $A$ be an irreducible of $\mathcal{P}_{\mathrm{fin},\times}(H)$, and suppose that $1_{H}\in uAv$ for some $u,v\in H^{\times}$. If $H$ is Dedekind-finite, then $uAv$ is irreducible in $\mathcal{P}_{\mathrm{fin},1}(H)$.
Proof.
If $|A|=2$, then $|uAv|=2$, and we know from Proposition 4.6 that every two-element set in $\mathcal{P}_{\mathrm{fin},1}(H)$ is irreducible.
Assume $|A|\geq 3$, and suppose that $H$ is Dedekind-finite. Then, by Proposition 4.10, AA is an atom of $\mathcal{P}_{\mathrm{fin},\times}(H)$, and hence, by [5, Proposition 3.1(iii)], the same is true of $uAv$. Since $1_{H}\in uAv$ (by hypothesis), it follows that $uAv$ is also an atom (and therefore an irreducible) of $\mathcal{P}_{\mathrm{fin},1}(H)$. ∎
Proposition 4.12.
If $H$ is Dedekind-finite, then $uAv$ is an irreducible of $\mathcal{P}_{\mathrm{fin},\times}(H)$ for all $u,v\in H^{\times}$ and every irreducible $A$ of $\mathcal{P}_{\mathrm{fin},1}(H)$.
Proof.
Let $A$ be an irreducible of $\mathcal{P}_{\mathrm{fin},1}(H)$, and suppose that $uAv=XY$ for some $u,v\in H^{\times}$ and $X,Y\in\mathcal{P}_{\mathrm{fin},\times}(H)$. We claim prove that
$uAv\text{ divides either }X\text{ or }Y\text{ in }\mathcal{P}_{\mathrm{fin},\times}(H).$
To begin, $A=\allowbreak u^{-1}XYv^{-1}$, and thus $1_{H}=u^{-1}xyv^{-1}$ for some $x\in X$ and $y\in Y$. Since $H$ is Dedekind-finite, $q:=u^{-1}x\in u^{-1}X$ is a unit of $H$, with inverse $q^{-1}=\allowbreak yv^{-1}\in Yv^{-1}$. Consequently, we can write $A=X^{\prime}Y^{\prime}$, where
$X^{\prime}:=u^{-1}Xq^{-1}\in\mathcal{P}_{\mathrm{fin},1}(H)\quad\text{and}\quad Y^{\prime}:=qYv^{-1}\in\mathcal{P}_{\mathrm{fin},1}(H).$
By the hypothesis that $A$ is irreducible in $\mathcal{P}_{\mathrm{fin},1}(H)$, either $X^{\prime}$ or $Y^{\prime}$ is divided by $A$ in the same monoid. Without loss of generality, suppose we are in the former case. There then exist $S,\allowbreak T\in\allowbreak\mathcal{P}_{\mathrm{fin},1}(H)$ such that $SAT=X^{\prime}=u^{-1}Xq^{-1}$, and hence
$X=(uSu^{-1})(uAv)(v^{-1}Tq).$
This shows that $uAv$ divides $X$ in $\mathcal{P}_{\mathrm{fin},\times}(H)$, as desired. Since $|uAv|=|A|\geq 2$, it follows from Theorem 4.2(i) that $uAv$ is an irreducible element of $\mathcal{P}_{\mathrm{fin},\times}(H)$. ∎
Corollary 4.13.
If $H$ is Dedekind-finite, every irreducible of $\mathcal{P}_{\mathrm{fin},\times}(H)$ is a quark.
Proof.
This is a direct consequence of Propositions 3.1, 4.6, and 4.10. ∎
Finale with lengths.
Another reason why most work on the arithmetic of (finitary) power monoids has focused so far on $\mathcal{P}_{\mathrm{fin},1}(H)$ is that, when $H$ is Dedekind-finite, $\mathcal{P}_{\mathrm{fin},1}(H)$ shares many arithmetic features with $\mathcal{P}_{\mathrm{fin},\times}(H)$ and, by Theorem 4.2(ii), the latter is divisor-closed in $\mathcal{P}_{\mathrm{fin}}(H)$. Details are worked out below.
Theorem 4.14.
If $H$ is Dedekind-finite, then $\mathcal{P}_{\mathrm{fin},1}(H)$ has the same system of lengths as $\mathcal{P}_{\mathrm{fin},\times}(H)$.
Proof.
Fix a non-unit $X\in\mathcal{P}_{\mathrm{fin},\times}(H)$, pick $u\in X\cap H^{\times}$, and set $X^{\prime}:=Xu^{-1}$. By Theorem 4.2(i), $\{1_{H}\}\neq X^{\prime}\in\mathcal{P}_{\mathrm{fin},1}(H)$. Let $L$ be the length set of $X$ in $\mathcal{P}_{\mathrm{fin},\times}(H)$, and $L^{\prime}$ be the length set of $X^{\prime}$ in $\mathcal{P}_{\mathrm{fin},1}(H)$. We claim that $L=L^{\prime}$.
To begin, let $A_{1}\ast\cdots\ast A_{n}$ be a length-$n$ factorization of $X$ in $\mathcal{P}_{\mathrm{fin},\times}(H)$. Then there exist $u_{1}\in A_{1}\cap H^{\times}$, …, $u_{n}\in A_{n}\cap H^{\times}$ such that $u=u_{1}\cdots u_{n}$. For every $i\in\allowbreak\llbracket 1,n\rrbracket$, define $v_{i}:=u_{1}\cdots u_{i}$ and $B_{i}:=v_{i-1}A_{i}v_{i}^{-1}$, where $v_{0}:=1_{H}$. We have from Proposition 4.11 that each $B_{i}$ is an irreducible of $\mathcal{P}_{\mathrm{fin},1}(H)$; in particular, note that $1_{H}=v_{i-1}u_{i}v_{i}^{-1}\in B_{i}$. Since $B_{1}\cdots B_{n}=v_{0}A_{1}\cdots A_{n}u^{-1}=A$, it follows that $B_{1}\ast\allowbreak\cdots\ast B_{n}$ is a length-$n$ factorization of $X^{\prime}$ in $\mathcal{P}_{\mathrm{fin},1}(H)$, thereby yielding $L\subseteq L^{\prime}$.
The reverse inclusion $L^{\prime}\subseteq L$ is immediate: if $C_{1}\ast\cdots\ast C_{m}$ is a length-$m$ factorization of $X^{\prime}$ in $\mathcal{P}_{\mathrm{fin},1}(H)$, then Proposition 4.12 implies that $C_{1}\ast\cdots\ast C_{m-1}\ast(C_{m}u)$ is a length-$m$ factorization of $X$ in $\mathcal{P}_{\mathrm{fin},\times}(H)$.
All in all, we see from the above that every length set of $\mathcal{P}_{\mathrm{fin},\times}(H)$ is also a length set of $\mathcal{P}_{\mathrm{fin},1}(H)$. Moreover, applying the same argument with $u=1_{H}$ under the additional assumption that $1_{H}\in X$, we obtain that every length set of $\mathcal{P}_{\mathrm{fin},1}(H)$ occurs in $\mathcal{P}_{\mathrm{fin},\times}(H)$. Thus, $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},\times}(H)$ have the same systems of lengths. ∎
With this point clarified, we know from Theorem 4.7 that every $X\in\mathcal{P}_{\mathrm{fin},1}(H)$ has a factorization (into irreducibles) regardless of any hypothesis on $H$. It is natural to ask whether any more refined conclusions can be drawn about these factorizations.
Theorem 4.15.
The following conditions are equivalent:
  1. (a)
    $H$ is torsion-free.
  2. (b)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is an FF monoid.
  3. (c)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is a BF monoid.
Proof.
The implication (a) $\Rightarrow$ (c) is straightforward from [5, Theorem 3.11(i)] and Corollary 4.8. As for the dual implication (c) $\Rightarrow$ (a), it is enough to consider that if $H$ has a non-identity element $x$ of finite order $n$, then
$X:=\{1_{H},x\}^{n}=\{1_{H},x\}^{n+k},\qquad\text{for all }k\in\mathbb{N}.$
By Proposition 4.6, this shows that the (infinite) interval $\llbracket n,\infty\rrbracket$ is contained in the length set of $X$, and hence $\mathcal{P}_{\mathrm{fin},1}(H)$ is not BF.
The equivalence (b) $\Leftrightarrow$ (c) is, on the other hand, a very special case of [10, Theorem 4.11(i)]. It suffices to note that, by Eq. (2), every $X\in\mathcal{P}_{\mathrm{fin},1}(H)$ has only finitely many divisors, and hence $\mathcal{P}_{\mathrm{fin},1}(H)$ is locally of finite type in the sense of [10, Definition 4.1(3)]. For additional details, see [10, Example 4.5(3)]. ∎
Corollary 4.16.
The following conditions are equivalent for a monoid $H$:
  1. (a)
    $H$ is trivial.
  2. (b)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is a UF monoid.
  3. (c)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is an HF monoid.
Proof.
We show that (c) $\Rightarrow$ (a), as the implications (a) $\Rightarrow$ (b) $\Rightarrow$ (c) are obvious.
Assume that $\mathcal{P}_{\mathrm{fin},1}(H)$ is HF. Since every HF monoid is BF, we have from Theorem 4.15 that $H$ is torsion-free. Suppose, for a contradiction, that $H$ is non-trivial, and let $x$ be a non-identity element of $H$. Then, by Proposition 4.6, the length set of
$X:=\{1_{H},x\}^{3}=\{1_{H},x\}\{1_{H},x^{2}\}$
is not a singleton. This, however, contradicts the hypothesis that $\mathcal{P}_{\mathrm{fin},1}(H)$ is HF. ∎
We conclude with some results concerning minimal factorizations.
Theorem 4.17.
The following conditions hold:
  1. (i)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is UmF-atomic if and only if $H$ is trivial.
  2. (ii)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is HmF-atomic if and only if $H$ is trivial or a cyclic group of order $3$.
  3. (iii)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is FmF and BmF.
Proof.
For (i) and (ii), see Theorem 4.14 and Corollary 4.15 in [5], respectively. As for (i), the conclusion follows directly from [10, Theorem 4.11], in a manner similar to the proof of the equivalence (b) $\Leftrightarrow$ (c) in Theorem 4.15. ∎
For the next result, we recall that a semigroup $S$ is called breakable if $xy=x$ or $xy=y$ for all $x,y\in S$. This is a rather special class of semigroups, first introduced by Rédei [33, Section 27], with a rich structural theory.
Theorem 4.18.
If $H$ is commutative, then the following conditions are equivalent:
  1. (a)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is UmF.
  2. (b)
    $H\setminus H^{\times}$ is a breakable subsemigroup of $H$, the unit group of $H$ is either trivial or cyclic of order two, and $ux=x$ for all $u\in H^{\times}$ and $x\in H\setminus H^{\times}$.
Proof.
See Theorem 4.8 in [13]. ∎
To date, there is no satisfactory characterization of when $\mathcal{P}_{\mathrm{fin},1}(H)$ is UmF in the non-commutative case, and the situation is even more dire for HmFness. For partial progress, see Lemma 3.4(i), Theorem 3.6, and Proposition 4.5 in [13].
5 Into the future.
In this final section, we present a selection of open problems of an arithmetic flavor, which we hope will stimulate further research on power monoids (and, more generally, on power semigroups) in the years to come.
A monoid is torsion if each of its elements has finite order (note that a torsion-free monoid is not a monoid that is not torsion). We begin with the following:
Conjecture 5.1.
If $H$ is not a torsion monoid, then every non-empty subset of the set of integers larger than one can be realized as the length set of a set $X\in\mathcal{P}_{\mathrm{fin},1}(H)$.
The conjecture was first stated in [16, Sect. 5] for the special case $H=\mathbb{N}$, which, however, turns out to be equivalent to the general case. Indeed, any monoid $H$ containing an element of infinite order has a submonoid isomorphic to $\mathbb{N}$. By Proposition 3.2 and Theorem 4.1(ii), it follows that every length set of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is also a length set of $\mathcal{P}_{\mathrm{fin},0}(H)$.
Now, we gather from Theorems 4.7(ii) and 4.15 that the monoid $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is BF and its irreducibles are all atoms. Hence, the conjecture reduces to proving that the system of length sets of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ consists of the singletons $\{0\}$ and $\{1\}$, together with all non-empty finite subsets of the integers greater than one.
As for partial progress, it is known from [16, Propositions 4.8–4.10] that, for each $n\geq 2$, the singleton $\{n\}$, the interval $\llbracket 2,n\rrbracket$, and the set $\{2,n\}$ are length sets of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$. This remained the state of the art until recently, and some authors (including the present one) started to cast doubt on the conjecture. A breakthrough came in summer 2025, when Reinhart [34] proved that $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is fully elastic: for every rational $q\geq 1$, there exists a length set $L$ of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ such that $\max L=q\min L$.
We continue with a question reminiscent of the “Characterization Problem for Systems of Lengths” in the classical theory of factorizations [20, Section 6].
Question 5.2 (Characterization Problem for Power Monoids).
Given a class $\mathcal{O}$ of finite monoids (for instance, the class of finite groups), determine whether there exists an integer $n\geq 1$ such that $H$ is isomorphic to $K$ for all $H,K\in\mathcal{O}$ with the property that (i) the systems of [minimal] length sets of $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ coincide and (ii) at least one of $H$ and $K$ has order greater than or equal to $n$.
The question is presented here for the first time. The answer is negative for the class of all finite monoids. Indeed, isomorphic monoids have the same systems of [minimal] length sets, and we gather from [47, Example 1.2(3)] that, for every $n\geq 1$, there exist non-isomorphic monoids $H$ and $K$ with $|H|=|K|=n$ whose reduced finitary power monoids are isomorphic. This naturally links Question 5.2 to Question 1.2.
We conjecture that the answer to Question 5.2 is positive for finite groups, for which Question 1.2 is known to have a positive answer [48]. Partial results for finite cyclic groups and minimal factorizations can be found in [5, Section 5], and there are at least a couple of non-trivial observations that are perhaps worth recording here.
First, we have from Proposition 3.2 and Theorem 4.1(ii) that the system of [minimal] length sets of a monoid $H$ is contained in the system of [minimal] length sets of $\mathcal{P}_{\mathrm{fin},1}(K)$ for every monoid $K$ into which $H$ embeds.
Second, [5, Proposition 4.12(i)] shows that, for every monoid $H$, the largest minimal length $\mu_{H}(X)$ of a set $X$ in $\mathcal{P}_{\mathrm{fin},1}(H)$ is bounded above by |X|1|X|-1. When $H$ is finite, it follows that the system of minimal length sets of $\mathcal{P}_{\mathrm{fin},1}(H)$ is contained in the power set of the interval $\llbracket 0,|H|-1\rrbracket$. Moreover, $H$ itself is an element of $\mathcal{P}_{\mathrm{fin},1}(H)$, and it seems reasonable to conjecture that $\mu_{H}(H)\to\infty$ as $|H|\to\infty$.
If confirmed, this latter conjecture would imply that, up to isomorphism, there are only finitely many pairs $(H,K)$ of finite monoids such that the systems of minimal lengths of $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ coincide.
Question 5.3.
Given a monoid $H$ and an integer $k\geq 1$, denote by $\mathscr{U}_{k}(\mathcal{P}_{\mathrm{fin},1}(H))$ and $\mathscr{U}_{k}^{\sf m}(\mathcal{P}_{\mathrm{fin},1}(H))$ the union of all length sets and the union of all minimal length sets of $\mathcal{P}_{\mathrm{fin},1}(H)$ containing $k$, respectively. Is $\mathscr{U}_{k}(\mathcal{P}_{\mathrm{fin},1}(H))$ an interval for every [large] $k\in\mathbb{N}^{+}$? What about $\mathscr{U}_{k}^{\sf m}(\mathcal{P}_{\mathrm{fin},1}(H))$? Note that the empty set is an interval.
Let $\mathbb{Z}_{n}$ be the additive group of integers modulo $n$, where $n\geq 2$. In [12, Example 3.12(2)], Cossu and Tringali noted that
[…]
and
$\mathscr{U}_{k}^{\sf m}(\mathcal{P}_{\mathrm{fin},0}(\mathbb{Z}_{n}))=\left\{\begin{array}[]{ll}\llbracket 2,n-1\rrbracket&\text{if }2\leq k<n,\\ \emptyset&\text{if }n\leq k.\end{array}\right$.
In addition, Theorem 4.7(ii) implies that $\mathscr{U}_{1}^{\sf m}(\mathcal{P}_{\mathrm{fin},0}(\mathbb{Z}_{n}))=\{1\}$.
On the other hand, we have already observed in the above that if $H$ contains an element of infinite order, then $\mathcal{P}_{\mathrm{fin},1}(H)$ has a divisor-closed submonoid isomorphic to $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$. By [16, Propositions 4.8], this implies that
$\llbracket 2,\infty\rrbracket\subseteq\mathscr{U}_{k}(\mathcal{P}_{\mathrm{fin},1}(H)),\qquad\text{for every integer }k\geq 2.$
One of the most intriguing aspects of power semigroups arises from the natural interplay between arithmetic and combinatorial properties. As a concrete example of this phenomenon (beyond the conjectures of Sárközy and Ostmann already mentioned in Section 1), let $H$ be a numerical monoid, and for all $k,l\in\mathbb{N}$ let $\alpha_{k,l}(H)$ denote the number of $k$-element atoms of $\mathcal{P}_{\mathrm{fin},0}(H)$ whose maximum is smaller than or equal to $l$. Experiments appear to support the following:
Conjecture 5.4 (Unimodality Conjecture).
For each $l\in\mathbb{N}$, there exists an index $k\in\allowbreak\llbracket 0,l\rrbracket$ such that $\alpha_{0,l}\leq\cdots\leq\alpha_{k,l}$ and $\alpha_{k,l}\geq\cdots\geq\alpha_{l,l}$.
The conjecture first appeared in an email from the author to Geroldinger and Gotti on January 20, 2023, where it was stated only for $H=\mathbb{N}$. It was later included as Problem 1 in the problem set of the “CrowdMath 2023: Arithmetic of Power Monoids” online collaborative project, organized by the same Gotti throughout 2023 and hosted on the Art of Problem Solving (AoPS) portal at
https://artofproblemsolving.com/polymath/mitprimes2023
As a first non-trivial step, Aggarwal et al. [1, Corollary 5.3] used a probabilistic approach to show that, in the special case $H=\mathbb{N}$, Conjecture 5.4 holds for almost all $l\in\mathbb{N}$ (here, “almost all” is understood in the sense of asymptotic density).
6 Acknowledgments.
The author was supported by the Natural Science Foundation of Hebei Province through grant A2023205045.
He thanks Federico Campanini (UCLouvain, Belgium), Laura Cossu (University of Cagliari, Italy), and Alfred Geroldinger (UniGraz, Austria) for numerous comments that have significantly improved the exposition. He is particularly grateful to Balint Rago (UniGraz, Austria) for a thorough proofreading of an earlier version.